# Integrating Large Language Models into Recommendation via Mutual Augmentation and Adaptive Aggregation

### v1

@@ 4.2.1. Data Augmentation for Direct Recommendation
For direct recommendation, the Bayesian Personalized Ranking (BPR) loss is commonly used to optimize the model (Rendle et al., 2012). The objective of BPR is to maximize the score difference between correctly recommended items and incorrectly recommended items, thereby improving the accuracy of recommendations. The BPR loss is defined as:
(2) $\mathcal{L}_{BPR}=-\sum_{(u,i,j)\in\mathcal{D}}\log\sigma(\hat{y}_{ui}-\hat{y}_{uj}),$
@@
where $(u,i,j)$ refers to a triple of user-item pairs, and the user $u$ has interacted with item $i$ (positive item) and item $j$ (negative item). $\mathcal{D}$ represents the set of such user-item pairs in the training data. $\hat{y}_{ui}$ denotes the predicted score or preference of user $u$ for item $i$.
@@
(3) $\mathcal{L}_{BPR}^{\prime}=-\sum_{(u,i,j)\in\mathcal{D}^{\prime}}\log\sigma(\hat{y}_{ui}-\hat{y}_{uj}).$
This data augmentation strategy leverages the power of the instruction tuned LLM to enhance the performance of the direct recommendation model.
4.2.2. Data Augmentation for Sequential Recommendation
[…]
4.2.3. Data Augmentation for Rating Prediction
In rating prediction tasks, we introduce the use of in-context learning (ICL) in LLMs to provide side information. This is primarily due to the fact that recommendation datasets may contain incomplete information. For instance, the popular Movielens dataset (Harper and Konstan, 2015) lacks information about the director of the movies, which can hinder the performance of a conventional rating prediction model. To mitigate this issue, we leverage the extensive world knowledge contained in an LLM. We prompt the LLM to provide side information, acting as additional attributes for users/items.
[…]
Top-k Recommendation Prompt Example:
Instruction: Rank the candidate movies based on user historical interactions and make the top k recommendations.
Interaction History: Beyond Rangoon (1995); Alien (1979); Hollow Reed (1996); Primary Colors (1998); …; Birds, The (1963)
Candidate Items: Last Dance (1996); Remains of the Day, The (1993); Assassins (1995); …; Fatal Instinct (1993)
Similar User Interaction History: L.A. Confidential (1997); Apt Pupil (1998); Kolya (1996); …; Star Wars (1977)
Conventional Model Prediction: Remains of the Day, The (1993); Addiction, The (1995); …; Fugitive, The (1993)
Output: Fugitive, The (1993); Angel Baby (1995); …; Remains of the Day, The (1993)
Rating Prediction Prompt Example:
Instruction: Predict the rating of a target movie based on the user’s historical movie ratings.
Rating History: Independence Day (1996): 3; Grosse Fatigue (1994): 3; Face/Off (1997): 4; …; Shall We Dance? (1996): 3
Candidate Item: Pink Floyd - The Wall (1982)
Similar User Rating History: L.A. Confidential (1997): 3; Apt Pupil (1998): 4; …; English Patient, The (1996): 3
Conventional Model Prediction: 3.2
Output: 3
Figure 2. Examples of instructions for top-k recommendation and rating prediction. The prompt augmentation component is underlined. To improve readability and facilitate better experimental evaluation, we introduce certain modifications to the original instructions employed in our experiments.
4.3. Prompt Augmentation for Large Language Model
Previous works (Bao et al., 2023; Zhang et al., 2023c) instruction tuning LLM for recommendation in a standard manner. However, these methods can be sub-optimal due to the challenges of distinguishing users based solely on text-based prompt descriptions. Although some concurrent studies (Zhang et al., 2023b; Zheng et al., 2023) incorporate unique identifiers to differentiate users, these approaches require complex semantic understanding of IDs and additional training, limiting their generalizability.
In this section, we introduce two text-based prompt augmentation strategies for LLM-based recommendations, i.e., we incorporate additional information within the prompt to enhance the model performance. First, we propose prompt augmentation with similar users, identifying users with analogous preferences to enrich the prompt, thereby enhancing the LLM’s ability to leverage collaborative information and generate personalized recommendations. Second, we propose prompt augmentation with conventional model prediction, providing prior knowledge to guide the LLM toward recommendations that align with user preferences. Collectively, these strategies harness the strengths of both LLMs and conventional recommendation models, ensuring generalizability across a wide range of recommendation scenarios. The illustration of prompt augmentation is underlined in Figure 2.
4.3.1. Prompt Augmentation with Collaborative Information from Similar User
To incorporate collaborative information within the prompt and facilitate LLM reasoning, we introduce a prompt augmentation strategy with similar user. Initially, we utilize a pre-trained conventional recommendation model to acquire embeddings for each user. These embeddings represent users in a latent space, which encapsulates their preferences and behaviors. Specifically, for a user $u$, in conjunction with a conventional recommendation model $\mathcal{M}_{c}$, we use $\mathcal{M}_{c}$ to obtain embeddings for each user, denoted as $\{e_{1},...,e_{n}\}$. These embeddings encapsulate the preferences and behaviors of the users, serving as a compact representation of the users in a latent space. We then calculate the similarity between these embeddings in the latent space. Various measurements, such as cosine similarity, Jaccard similarity, and Euclidean distance, could be employed in this context. In this paper, we calculate the cosine similarity to measure how closely two vectors align, denoted as:
(4) $sim(u,v)=\frac{e_{u}\cdot e_{v}}{||e_{u}||\cdot||e_{v}||},$
@@ 4.3.1. Prompt Augmentation with Collaborative Information from Similar User
where $e_{u}$ and $e_{v}$ are the embeddings of user $u$ and $v$, respectively, and $u,v\in\mathcal{U}$. The |||||$\cdot$||| | ⋅ | | denotes the Euclidean norm and $\cdot$ denotes the dot product. We identify the pair of users $(u,v)$ that have the highest similarity, indicating that they are the most similar in terms of their preferences and behaviors. We then use the items interacted with by the most similar user to enrich the prompt for the target user. This strategy leverages the collaborative information gleaned from similar users to generate more relevant and accurate prompts, thereby enhancing the recommendation performance of the LLM.
@@
4.3.2. Prompt Augmentation with Prior Knowledge from Conventional Recommendation Model Prediction
@@
Notably, unlike ID-based methods such as (Zhang et al., 2023b; Zheng et al., 2023), our approach relies entirely on text, enabling easy adaptation to new situations. Also, the prompt augmentation could be used as a plug-and-play component for recommendation with closed source LLM, such as the GPT-4 model (OpenAI, 2023).
4.4. Adaptive Aggregation
We endeavor to aggregate the outputs of LLM and conventional recommendation models at the result level for improved performance, considering the disparate model structures. However, indiscriminate aggregation of model predictions can potentially lead to suboptimal results. Conventional recommendation models, known for their susceptibility to the long tail issue, often struggle when dealing with the tail segment. In contrast, LLMs, by leveraging contextual information, are able to maintain a relatively uniform performance across all segments. Motivated by these observations, we first define the long-tail coefficient and subsequently adaptively aggregate the predictions from both model types.
@@ 4.4. Adaptive Aggregation
(5) $\ell_{u}=\text{log}\left(N({u})+1\right),$
where $N({u})$ is the number of interaction for user $u$. A lower long-tail coefficient value indicates that the user has fewer feedback.
@@
4.4.1. Adaptive Aggregation for Rating Prediction
For the rating prediction task, we employ an instruction-tuned LLM to predict user-item utility scores directly. This approach incorporates the understanding of complex semantics and context by the LLM, which might be overlooked by traditional models. Similarly, conventional recommendation methods leverage collaborative information and user/item features for predicting the user rating. Specifically, the utility weight for user $u$, denoted as $U_{u}$, is directly set as its user rating. Subsequently, the LLM is engaged to predict the rating, symbolized as $U_{LLM}$. There are various methods to derive a final result based on the utility scores, such as training a neural network to process the utility scores from LLM and conventional models, yielding a final output via learning the complex reflection. However, for the sake of simplicity in this paper, we adopt a simple yet effective linear interpolation approach. The final utility score for a user $u$ amalgamates the values from both models, represented as:
(6) $U_{u}=\alpha_{u}{U}_{LLM}+(1-\alpha_{u})U_{Rec},$
@@ 4.4.1. Adaptive Aggregation for Rating Prediction
(7) $\alpha_{u}=\text{max}\left(\frac{\ell_{max}-\ell_{u}}{\ell_{max}-\ell_{min}},\alpha_{2}\right)\cdot\alpha_{1},$
@@
4.4.2. Adaptive Aggregation for Top-$k$ Recommendation
@@
For the top-$k$ recommendation task, the LLM is employed to re-rank the item list generated by a conventional recommendation model. Specifically, from conventional recommendation methods, we curate a top-ranked list comprising $k^{\prime}$ items, denoted as $\{i_{1},...,i_{k^{\prime}}\}$. Each item in this list is assigned a utility weight, ${U}^{i}_{Rec}=-s\cdot\mathcal{C}$, where $\mathcal{C}$ is a constant and $s$ represents the position of item $i$, i.e., $s\in\{1,...,k^{\prime}\}$. A higher utility weight indicates a stronger inclination of the user’s preference. For listwise comparison conducted by the LLM, the process begins by using the LLM to directly output the predicted order of these candidate items. Then we assign utility scores for items at each position, denoted as $U_{1},{U}_{2},...,U_{k^{\prime}}$, where $U_{1}\geq U_{2}\geq...\geq U_{k^{\prime}}$. The final utility score for an item amalgamates the values from both the original rating and the LLM prediction, similar to the Equation (6).
4.5. Training Strategy for LLM
4.5.1. Instruction Tuning Dataset Construction
@@
This section details the creation of an instruction-tuning dataset that encompasses two types of recommendation tasks catering to top-$k$ recommendation and rating prediction scenarios. A depiction of these two tasks, specifically referred to as listwise ranking and rating prediction, can be found in Figure 2. It is noteworthy that we also employ the LLM to execute pointwise ranking within top-$k$ recommendation scenarios, i.e., utilizing LLM to predict ratings for each item within the top-k recommendations and sorting the predicted ratings to derive the final result.
4.5.2. Optimization via Instruction Tuning
@@
In this work, we perform full parameter instruction tuning to optimize LLMs using generated instruction data. Due to our need for customization, we chose LLaMA-2 (Touvron et al., 2023b), an open-source, high-performing LLM, which permits task-specific fine-tuning. During supervised fine-tuning, we apply a standard cross-entropy loss following Alpaca (Taori et al., 2023). The training set $\mathcal{D}_{ins}$ consists of instruction input-output pairs $(x,y)$, which have been represented in natural language. The objective is to fine-tune the pre-trained LLM by minimizing the cross-entropy loss, formalized as:
(8) $\min_{\Theta}\sum_{(x,y)\in\mathcal{D}_{ins}}\sum_{t=1}^{|y|}-\log P_{\Theta}\left(y_{t}\mid x,y_{[1:t-1]}\right),$
@@
where $\Theta$ are the original parameters for LLM, $P_{\Theta}$ is the conditional probability, $|y|$ is the number of tokens in $y$, $y_{t}$ is the $t$-th token in the target output $y$, and $y_{[1:t-1]}$ represents tokens preceding $y_{t}$ in $y$. By minimizing this loss function, the model fine-tunes its parameters $\Theta$ to adapt to the specifics of the new instruction tuning dataset $\mathcal{D}_{ins}$, while leveraging the general language understanding and reasoning that has been acquired during pre-training (Zhang et al., 2023a). In this manner, LLM can capture the user’s preferences for items expressed in natural language, facilitating diverse recommendation tasks, including top-k recommendation and rating prediction.
5. Experiment

### v2

@@ III-B1 Data Augmentation for Direct Recommendation
For direct recommendation, the Bayesian Personalized Ranking (BPR) loss [46] is commonly used to optimize the model. The objective of BPR is to maximize the score difference between correctly recommended items and incorrectly recommended items, thereby improving the accuracy of recommendations. The BPR loss is defined as:
$\mathcal{L}_{BPR}=-\sum_{(u,i,j)\in\mathcal{D}}\log\sigma(\hat{y}_{ui}-\hat{y}_{uj}),$ (1)
@@
where $(u,i,j)$ refers to a triple of user-item pairs, and the user $u$ has interacted with item $i$ (positive item) and item $j$ (negative item). $\mathcal{D}$ represents the set of such user-item pairs in the training data. $\hat{y}_{ui}$ denotes the predicted score or preference of user $u$ for item $i$. $\sigma$ is the sigmoid function and $\sigma(\hat{y}{ui}-\hat{y}{uj})$
Inspired by this, we propose a data augmentation strategy that harnesses the capabilities of the LLM to enhance the direct recommendation. Initially, we randomly select pairs of items for a user $u$ and prompt the LLM to rank each pair based on the user’s likely preference. The ranking prediction based on LLM is then combined with the original data and used to train a direct recommendation model. Formally, let $(i_{u1},i_{u2})$ denote a pair of un-interacted items for a user $u$. The LLM is prompted to rank these items, denoted as $i^{+},i^{-}=\textit{LLM}(\mathcal{P}_{1})$, where $\mathcal{P}_{1}$ is the corresponding prompt and $i^{+}$ is the item preferred over $i^{-}$. The training data $\mathcal{D}$ is updated as $\mathcal{D}^{\prime}=\mathcal{D}\cup(u,i^{+},i^{-})$. The BPR loss is then updated as:
$\mathcal{L}_{BPR}^{\prime}=-\sum_{(u,i,j)\in\mathcal{D}^{\prime}}\log\sigma(\hat{y}_{ui}-\hat{y}_{uj}).$ (2)
III-B2 Data Augmentation for Sequential Recommendation
For sequential recommendation, the data augmentation strategy involves enriching the sequence of interacted items with additional items predicted by the LLM. Let’s consider a user $u$ with a corresponding sequence of interacted items $\{i_{1},...,i_{l}\}$. We randomly sample a list of un-interacted items $\{i_{u1},...,i_{ut}\}$, and adopt the prompt $\mathcal{P}_{2}$ to ask the LLM to predict the item most likely to be preferred by the user, denoted as $i_{p}=\textit{LLM}(\mathcal{P}_{2})$. This predicted item $i_{p}$ is then randomly inserted into the user’s sequence, resulting in an augmented sequence $\{i_{1},...,i_{p},...,i_{l}\}$. This sequence is used to train a more powerful sequential recommendation model.
III-B3 Data Augmentation for Rating Prediction
In rating prediction tasks, we introduce the use of in-context learning (ICL) in LLMs to provide side information. This is primarily due to the fact that recommendation datasets may contain incomplete information. For instance, the popular Movielens dataset [13] lacks information about the director of the movies, which can hinder the performance of a conventional rating prediction model. To mitigate this issue, we leverage the extensive world knowledge contained in an LLM. We prompt the LLM to provide side information, acting as additional attributes for users/items.
Formally, we denote the rating prediction model as $\mathcal{M}_{r}$, and the attribute set as $\mathcal{A}=\{a_{1},a_{2},...,a_{n}\}$, where $a_{i}\in\mathcal{A}$ denotes a distinct attribute. The model predicts the rating as $\mathtt{Pred}=\mathcal{M}_{r}(\mathcal{A})$. We then prompt the LLM to provide additional attributes, where the prompt $\mathcal{P}_{3}$ contains some corresponding examples followed by detailed instructions. The process is denoted as $\{a_{d1},a_{d2},...\}=\textit{LLM}(\mathcal{P}_{3})$. The augmented attribute set is then formed as $\mathcal{A}^{\prime}=\mathcal{A}\cup\{a_{d1},a_{d2},$...\}. The model then predicts the rating using the augmented attribute set, denoted as $\mathtt{Pred}^{\prime}=\mathcal{M}_{r}(\mathcal{A}^{\prime})$. This approach allows us to leverage the LLM’s world knowledge to enhance the performance of the rating prediction model.
III-C Prompt Augmentation for LLM
Previous works [3, 66] instruction tuning LLM for recommendation in a standard manner. However, these methods can be sub-optimal due to the challenges of distinguishing users based solely on text-based prompt descriptions. Although some concurrent studies [69, 72] incorporate unique identifiers to differentiate users, these approaches require complex semantic understanding of IDs and additional training, limiting their generalizability.
We introduce two text-based prompt augmentation strategies for LLM-based recommendations, i.e., we incorporate additional information within the prompt to enhance the model performance. First, we propose prompt augmentation with similar users, identifying users with analogous preferences to enrich the prompt, thereby enhancing the LLM’s ability to leverage collaborative information and generate personalized recommendations. Second, we propose prompt augmentation with conventional recommendation model prediction, providing prior knowledge to guide the LLM toward recommendations that align with user preferences. Collectively, these strategies harness the strengths of both LLMs and conventional recommendation models, ensuring generalizability across a wide range of recommendation scenarios. The illustration of prompt augmentation is underlined in Figure 3.
III-C1 Prompt Augmentation with Collaborative Information from Similar Users
To incorporate collaborative information within the prompt and facilitate LLM reasoning, we introduce a prompt augmentation strategy with similar users. Initially, we utilize a pre-trained conventional recommendation model to acquire embeddings for each user. These embeddings serve as a representation of users in a latent space, which encapsulates their preferences and behaviors. Specifically, for a user $u$, in conjunction with a conventional recommendation model $\mathcal{M}_{c}$, we use $\mathcal{M}_{c}$ to obtain embeddings for each user, denoted as $\{e_{1},...,e_{n}\}$. We then calculate the similarity between these embeddings in the latent space. Various measurements, such as cosine similarity, Jaccard similarity, and Euclidean distance, could be employed in this context. In this paper, we calculate the cosine similarity to measure how closely two vectors align, denoted as:
@@ III-C1 Prompt Augmentation with Collaborative Information from Similar Users
where $e_{u}$ and $e_{v}$ are the embeddings of user $u$ and $v$, respectively, and $u,v\in\mathcal{U}$. The $||\cdot||$ denotes the Euclidean norm and $\cdot$ denotes the dot product. We identify the pair of users $(u,v)$ that have the highest similarity, indicating that they are the most similar in terms of their preferences and behaviors. We then use the items interacted with by the most similar user to enrich the prompt for the target user. This strategy leverages the collaborative information gleaned from similar users to generate more relevant and accurate prompts, thereby enhancing the recommendation performance of the LLM.
@@
III-C2 Prompt Augmentation with Prior Knowledge from Conventional Recommendation Model Prediction
@@
Notably, unlike ID-based methods such as [69, 72], our approach relies entirely on text, enabling easy adaptation to new situations. Moreover, the prompt augmentation could be used as a plug-and-play component for recommendation with closed source LLM, such as the GPT-4 model [41].
III-D Adaptive Aggregation
We endeavor to merge the textual and collaborative modality of both the LLM and conventional recommendation models at the post-learning stage. Considering the disparate model structures, we aggregate the outputs from both types of models for improving recommendation performance. However, indiscriminate aggregation of model predictions can lead to sub-optimal results. Conventional recommendation models, known for their susceptibility to the long tail issue, often struggle when dealing with the tail segment. In contrast, LLMs, by leveraging contextual information, are able to maintain a relatively uniform performance across all segments. Motivated by these observations, we first define the long-tail coefficient and subsequently adaptively aggregate the predictions from both model types.
@@ III-D Adaptive Aggregation
$\ell_{u}=\text{log}\left(N({u})+1\right),$ (4)
where N(u)N({u}) is the number of interaction for user $u$. A lower long-tail coefficient value indicates that the user has fewer feedback.
@@
III-D1 Adaptive Aggregation for Rating Prediction
For the rating prediction task, we employ an instruction-tuned LLM to predict user-item utility scores directly. This approach incorporates the understanding of complex semantics and context by the LLM, which might be overlooked by conventional models. Similarly, conventional recommendation methods leverage collaborative information and user/item features for predicting the user rating. Specifically, the LLM is adopted to predict the user rating as utility score, denoted as $U_{LLM}$. Besides, the conventional predicts the user rating as $U_{Rec}$. Subsequently, there are various methods to derive a final utility score based on both kinds of models, such as training a neural network to process the utility scores from LLM and conventional models, yielding a final output via learning a complex reflection. However, for the sake of simplicity in this paper, we adopt a simple yet effective linear interpolation approach. Thus, the final utility score for a user $u$ amalgamates the values from both models, represented as:
$U_{u}=\alpha_{u}{U}_{LLM}+(1-\alpha_{u})U_{Rec},$ (5)
@@ III-D1 Adaptive Aggregation for Rating Prediction
$\alpha_{u}=\text{max}\left(\frac{\ell_{max}-\ell_{u}}{\ell_{max}-\ell_{min}},\alpha_{2}\right)\cdot\alpha_{1},$ (6)
@@
III-D2 Adaptive Aggregation for Top-$k$ Recommendation
@@
For the top-$k$ recommendation task, the LLM is employed to re-rank the item list generated by a conventional recommendation model. Specifically, from conventional recommendation methods, we curate a top-ranked list comprising $k^{\prime}$ items, denoted as $\{i_{1},...,i_{k^{\prime}}\}$. Each item in this list is assigned a utility weight, ${U}^{i}_{Rec}=-s\cdot\mathcal{C}$, where $\mathcal{C}$ is a constant and $s$ represents the position of item $i$, i.e., $s\in\{1,...,k^{\prime}\}$. A higher utility weight indicates a stronger inclination of the user’s preference. For listwise comparison conducted by the LLM, the process begins by using the LLM to directly output the predicted order of these candidate items. Then we assign utility scores for items at each position, denoted as $U_{1},{U}_{2},...,U_{k^{\prime}}$, where $U_{1}\geq U_{2}\geq...\geq U_{k^{\prime}}$. In this paper, we also set ${U}_{p}=-p\cdot\mathcal{C}$ where ${p}\in\{1,2,...,{k^{\prime}}\}$, and Up{U}{p} is the utility score assigned to the item at position $p$. The final utility score for an item amalgamates the values from both the original rating and the LLM prediction, similar to Equation (5).
III-E Training Strategy for LLM
III-E1 Instruction Tuning Dataset Construction
@@
This section details the creation of an instruction-tuning dataset that encompasses two types of recommendation tasks catering to top-$k$ recommendation and rating prediction scenarios. A depiction of these two tasks, specifically referred to as listwise ranking and rating prediction, can be found in the Figure 3. It is noteworthy that we also employ the LLM to execute pointwise ranking within top-$k$ recommendation scenarios, i.e., utilizing LLM to predict ratings for each item within the top-k recommendations and sorting the predicted ratings to derive the final result. In our dataset, the input $x$ consists of user and item features (e.g., user preferences, item attributes) used to prompt the model, while the output $y$ represents the corresponding recommendation or prediction. The ground-truth information from the training set is used to construct the output $y$, ensuring that the model learns from accurate and verified data.
Figure 3: Examples of instructions for top-k recommendation and rating prediction. The prompt augmentation component is highlighted. The similar user histories and conventional model predictions are prompt augmented, which are highlighted in light purple and beige. We keep the prompt template structure and convert the bullet-point format into cohesive sentences for use in experiments. In the top-k recommendation, both direct and sequential recommendations follow the same prompt structure, with the key difference being that the user interaction history for sequential recommendations is time-series aware.
III-E2 Optimization via Instruction Tuning
@@
In this work, we perform full parameter instruction tuning to optimize LLMs using generated instruction data. Due to our need for customization, we chose LLaMA-2 [52], an open-source, high-performing LLM, which permits task-specific fine-tuning. During supervised fine-tuning, we apply a standard cross-entropy loss following Alpaca [50]. The training set $\mathcal{D}_{ins}$ consists of instruction input-output pairs $(x,y)$, which have been represented in natural language. The objective is to fine-tune the pre-trained LLM by minimizing the cross-entropy loss, formalized as:
$\min_{\Theta}\sum_{(x,y)\in\mathcal{D}_{ins}}\sum_{t=1}^{|y|}-\log P_{\Theta}\left(y_{t}\mid x,y_{[1:t-1]}\right),$ (7)
@@
where $\Theta$ are the original parameters for LLM, $P_{\Theta}$ is the conditional probability, $|y|$ is the number of tokens in $y$, $y_{t}$ is the $t$-th token in the target output $y$, and $y_{[1:t-1]}$ represents tokens preceding $y_{t}$ in $y$. By minimizing this loss function, the model fine-tunes its parameters $\Theta$ to adapt to the specifics of the new instruction tuning dataset $\mathcal{D}_{ins}$, while leveraging the general language understanding and reasoning that has been acquired during pre-training [67]. In this manner, LLM can capture the user’s preferences for items expressed in natural language, facilitating diverse recommendation tasks, including top-k recommendation and rating prediction.
TABLE II: Performance achieved by different direct recommendation methods. H@$k$ indicates the hit rate at position $k$, and N@$k$ represents normalized discounted cumulative gain (NDCG) at position $k$.
IV Experiment